# Reasoning on Time-Series for Financial Technical Analysis

### v1

@@
Abstract.
While Large Language Models have been used to produce interpretable stock forecasts, they mainly focus on analyzing textual reports but not historical price data, also known as Technical Analysis. This task is challenging as it switches between domains: the stock price inputs and outputs lie in the time-series domain, while the reasoning step should be in natural language. In this work, we introduce Verbal Technical Analysis (VTA), a novel framework that combine verbal and latent reasoning to produce stock time-series forecasts that are both accurate and interpretable. To reason over time-series, we convert stock price data into textual annotations and optimize the reasoning trace using an inverse Mean Squared Error (MSE) reward objective. To produce time-series outputs from textual reasoning, we condition the outputs of a time-series backbone model on the reasoning-based attributes. Experiments on stock datasets across U.S., Chinese, and European markets show that VTA achieves state-of-the-art forecasting accuracy, while the reasoning traces also perform well on evaluation by industry experts.
Large Language Models, Time-Series, Technical Analysis
copyright: rightsretainedjournalyear: 2025conference: Workshop on Rethinking Financial Time-Series (ICAIF ’25); Nov 16 2025; Singapore, Singaporeccs: Computing methodologies Temporal reasoningccs: Applied computing Forecastingccs: Applied computing Economics
Table 1. Comparison of relevant works. Our work contributes a novel explainable financial signal for practitioners and produces some insights into how time-series forecasting can be made interpretable.
@@ Abstract.
Fin-R1 (liu2025fin), Fino1 (qian2025fino1) Financial Question Answering Financial Reports Textual Answers
FinMem (yu2025finmem), FinCon (yu2024fincon) Investment Decision-Making Text, Tabular, Audio Binary (Buy/Sell)
GPT-4 (yu2023temporal), SEP (koa2024learning) Market Direction Forecasting News, Social Texts Binary or Quantile-Based
@@
Time-LLM (jin2023time), CALF (liu2025calf) Time-Series Forecasting General Time-Series Time-Series Only
@@
Verbal Technical Analysis (VTA) Financial Time-Series Forecasting Financial Time-Series Time-Series + Reasoning
1. Introduction
With the advent of Large Language Models (LLMs), an increasingly popular application is in financial analysis (wu2023bloomberggpt; xie2023pixiu). This spans a wide range of tasks, including financial question answering (FinQA) (liu2025fin; qian2025fino1), investment decision-making (yu2025finmem; yu2024fincon), and market forecasting (yu2023temporal; koa2024learning). Majority of existing approaches primarily utilize the strong natural-language capabilities of LLMs to analyze financial reports or do sentiment analysis on social texts (see Table 1), but neglects interpretable analysis on stock price data, which arguably contain useful information for financial practitioners. This is likely due to the fact that there are not many known techniques on time-series reasoning in the general domain (merrill2024language; chow2024towards). However, unlike other time-series data, financial time-series contains known interpretable signals which are widely studied by experts, known as Technical Analysis (kirkpatrick2010technical). We explore using these signals to verbally analyze financial time-series data, to produce stock forecasts with interpretable reasoning traces.
@@ 1. Introduction
The use of LLMs for time-series reasoning is hindered by three main challenges. Firstly, current LLMs have limited capabilities in time-series forecasts. Some works have tackled this by modifying the embedding space to produce time-series outputs (jin2023time; liu2025calf), but this comes at the cost of interpretability as the LLM loses its natural language capability. Secondly, on a higher level, current LLMs are not known to have the ability to do verbal reasoning on time-series to produce accurate forecasts (merrill2024language; chow2024towards). This involves understanding how to best analyze the predictive signals in the time-series data in an unsupervised manner. Thirdly, the reasoning trace of the LLM further needs to be converted into time-series output to produce useful stock forecasts. LLMs are typically fine-tuned on next-token predictions (radford2019language), and using direct time-series outputs would not produce the best forecasts, which we verify empirically.
@@
To address these problems, we present three key contributions. Firstly, we propose our Verbal Technical Analysis (VTA) framework, which combines a backbone time-series model (which we termed as “latent thinking”) with a reasoning LLM (termed as “verbal reasoning”) to produce interpretable stock time-series forecasts. This framework combines the strong pattern processing ability of state-of-the-art time-series models and the strong reasoning ability of LLMs to produce forecasts that are both accurate and interpretable. Secondly, for reasoning over time-series, the stock time-series data is converted into textual annotations (lin2024decoding) as inputs to the LLM. The reasoning trace is then optimized through a modified Group Relative Policy Optimization (GRPO) objective (shao2024deepseekmath) that uses an inverse Mean Squared Error (MSE) reward scoring, which we termed as Time-GRPO. Thirdly, to produce time-series outputs from the reasoning traces, we condition (ho2022classifier) the generated outputs from the time-series model on the reasoning-based attributes.
@@
To demonstrate the effectiveness of VTA, we perform extensive experiments across established stock baselines (xu2018stock) and additional stock data across the U.S., Chinese, and European markets. We show that our model forecasts achieve state-of-the-art results in prediction accuracy, while also being interpretable. In addition, evaluation by industry experts show that the reasoning traces score high across various evaluation metrics from literature. Finally, to justify the practical capability of the model, we form Markowitz portfolios across the prediction length and show that the portfolios formed from VTA forecasts also perform well on investment metrics.
2. Related Works
@@
Financial Large Language Models. The rise of Large Language Models (LLMs) has spurred a growing body of research on their application in finance. The earliest works focus on developing general-purpose financial LLMs, such as BloombergGPT (wu2023bloomberggpt) and FinMA (xie2023pixiu), by finetuning on a large set of financial corpora across multiple downstream tasks. Later works began to tackle the specific challenges of LLMs in finance. For example, works on financial question-answering (FinQA) (liu2025fin; qian2025fino1) focus on teaching LLMs to analyze financial reports, which requires the ability to read structured financial tables and extract insights from complex documents (zhu2021tat). LLMs for investment decision-making (yu2025finmem; yu2024fincon) typically utilize multi-agent systems to handle different parts of an investment decision, including but not limited to document analysis, memory, risk control, etc. LLMs for financial forecasting (yu2023temporal; koa2024learning) seeks to predict the direction in which the market will go. Typically, these works analyze textual sources in order to understand the sentiment or financial health of a company. Our work positions itself in this field by learning to produce interpretable forward-looking signals from financial time-series, which could benefit all applications above.
@@
Time-Series Large Language Models. At the same time, there is also a growing body of research in utilizing LLMs in the time-series domain. Early works (xue2023promptcast; gruver2023large) utilized the use of their sequence completion capabilities by reframing the time-series forecasting task, prompting the LLM to complete a given string of numbers. Later works (jin2023time; liu2025calf) leveraged on their large scale parameters and robust pattern recognition by fine-tuning LLMs for time-series forecasting tasks. However, these approaches typically modify the embedding space of the LLMs, making them lose their original language reasoning capabilities. Some works have also started to explore the ability of LLMs to reason over time-series (kong2025time). It was found that language models are “remarkably bad” at zero-shot time-series reasoning (merrill2024language), whereas fine-tuning them using a latent encoder (chow2024towards) shows some promising early results in reasoning over time-series for captioning (but not forecasting). Our work builds on this line of research by exploiting the predictive signals in financial time-series to produce interpretable forecasts, providing some insights on the capabilities of LLMs in reasoning over time-series.
3. Verbal Technical Analysis
@@
Figure 1. The Verbal Technical Analysis (VTA) framework. We first teach an LLM to reason over time-series data. The reasoning outputs are used to condition a time-series forecasting model, to produce forecasts with similar attributes. This results in forecasts with interpretable reasoning traces.
3.1. Problem Formulation
We consider the task of forecasting short-term stock prices, based on a historical window of $T$ days. Let $\mathbf{X}=\{\mathbf{x}_{t-T+1},\mathbf{x}_{t-T+2},\cdots,\mathbf{x}_{t}\}$, where the input vector consists of the open price, high price, low price, volume traded, closing price and adjusted closing price, i.e., $\mathbf{x}_{t}=\left[o_{t},h_{t},l_{t},v_{t},c_{t},p_{t}\right]$. We aim to generate an output $\mathbf{Y}=\{\mathbf{v},\mathbf{y}\}$, which consists of the verbal reasoning trace $\mathbf{v}$ and the price trajectory over the next $T^{\prime}$ trading days $\mathbf{y}=\{p_{t+1},p_{t+2},\cdots,p_{t+T^{\prime}}\}$.
3.2. Time-Series Reasoning
To teach an LLM to reason over time-series inputs, we use a textual annotator to extract useful interpretable signals for forecasting. The LLM uses these indicators to reason over the time-series to make forecasts without any supervision data. This is achieved through our proposed Time-Series Group Relative Policy Optimization (Time-GRPO) method, which uses a multi-stage reinforcement learning (RL) pipeline, together with a modified GRPO objective (shao2024deepseekmath).
The time-series input is first converted into textual annotations (lin2024decoding), which consist of its statistics information (jin2023time) (e.g., its mean, minimum and maximum values) and financial technical indicators (murphy1999technical) (e.g., moving averages, momentum, etc. ). Formally, we have:
(1) $\mathbf{X^{\prime}}=\mathbf{f}(\mathbf{X}),$
where $\mathbf{f}$ is the annotation function and $\mathbf{X^{\prime}}$ are the annotated values.
@@ 3.2. Time-Series Reasoning
(2) $\mathcal{L}_{\text{time-grpo}}(\theta)=\mathbb{E}_{\mathbf{q}\sim\mathcal{Q},\{\mathbf{o_{i}}\}_{i=1}^{G}\sim\pi_{\theta_{\text{old}}}(\mathbf{O}|\mathbf{q})}\Bigg[\frac{1}{G}\sum_{i=1}^{G}\Big(\min\Big(\frac{\pi_{\theta}(\mathbf{o_{i}}|\mathbf{q})}{\pi_{\theta_{\text{old}}}(\mathbf{o_{i}}|\mathbf{q})}A_{i},\\ \text{clip}\Big(\frac{\pi_{\theta}(\mathbf{o_{i}}|\mathbf{q})}{\pi_{\theta_{\text{old}}}(\mathbf{o_{i}}|\mathbf{q})},1-\epsilon,1+\epsilon\Big)A_{i}\Big)-\beta\,\mathbb{D}_{\text{KL}}(\pi_{\theta}\|\pi_{\text{ref}})\Big)\Bigg],\\ \!\!\!\!\!\!\!\!\mathbb{D}_{\text{KL}}(\pi_{\theta}\|\pi_{\text{ref}})=\frac{\pi_{\text{ref}}(\mathbf{o_{i}}|\mathbf{q})}{\pi_{\theta}(\mathbf{o_{i}}|\mathbf{q})}-\log\frac{\pi_{\text{ref}}(\mathbf{o_{i}}|\mathbf{q})}{\pi_{\theta}(\mathbf{o_{i}}|\mathbf{q})}-1,$
where $\epsilon$ and $\beta$ are hyper-parameters. $A_{i}$ denotes the advantage of the LLM policy, which is derived from a set of rewards $\{r_{1},r_{2},\cdots,r_{G}\}$ that are associated with outputs $\mathbf{O}$ in each group:
(3) $A_{i}=\frac{r_{i}-\text{mean}(\{r_{1},r_{2},\cdots,r_{G}\})}{\text{std}(\{r_{1},r_{2},\cdots,r_{G}\})}.$
We utilize the format reward that was used in previous works (guo2025deepseek), that enforces the model to always employ a thinking process that is between <think><think> and </think></think> tags.
Ideally, the generated reasoning trace should also maximize the expected accuracy of the time-series forecasts. This is achieved by using the Mean-Squared Error (MSE) score as an additional reward:
(3) $r_{\text{MSE}}(\theta)=1/\left(\lambda\cdot\left\|\hat{\mathbf{y}}_{\theta}-\mathbf{y}\right\|_{2}^{2}\right),$
@@
Training Pipeline. Following established practices in LLM fine-tuning literature (guo2025deepseek; ouyang2022training; lu2024deepseek; chow2024towards), Time-GRPO utilizes a multi-stage pipeline to fine-tune the time-series reasoning LLM.
The first stage represents the cold-start phase (guo2025deepseek). As there were no “gold-standard” supervision data, this stage is used for generating the initial training samples, guided by the $\mathcal{L}_{\text{time-grpo}}$ objective. Empirically, we find that the forecasting performance would not significantly improve here, but the process lets us generate training data for the next stage to fine-tune the base model.
The second stage focuses on teaching the model to produce more effective reasoning. This is achieved through rejection sampling, where we keep only the reasoning traces that lead to forecasts with lower Mean Squared Error (MSE). To ensure better consistency of training samples, we also bucket the samples across different stocks and time-periods, and filter for those with MSE in the bottom $10^{\text{th}}$ percentile in each bucket. The model is then trained on these filtered samples through the use of supervised fine-tuning (SFT).
The third stage optimizes the model for the best forecasting performance. Given that the model has learnt to reason over time-series data in the previous stage, this stage now aims to search for the best reasoning policy that can maximize the expected accuracy of the predicted time-series. For this stage, the model is also optimized using the $\mathcal{L}_{\text{time-grpo}}$ objective.
3.3. Time-Series Forecasting
To perform time-series forecasting, we employ LLM-based time-series models. Works have shown that the powerful contextual modeling capabilities of LLMs can be effectively adapted for time-series forecasting tasks (zhou2023one; chang2023llm4ts; cao2023tempo). A key technique is to align the time-series and language distributions (sun2023test; jin2023time) such that the model is able to understand the context of time-series data (e.g., up, down, steady, etc. ). For our backbone model, we repurpose an LLM for cross-modal fine-tuning (liu2025calf).
For this step, we first pass the time-series input $\mathbf{X}$ through an embedding layer, followed by a multi-head attention layer, to obtain the projected time tokens $\mathbf{X}_{\text{time}}$. Next, it is observed that similar words are usually close to each other in the LLM embedding space, and for non-text based tasks (e.g., time-series forecasting), it is sufficient to keep cluster centers to reduce training costs (sun2023test; liu2025calf). To do this, we perform Principal Component Analysis (PCA) to retrieve the principal word embeddings $\mathbf{\hat{D}}$. We then pass the projected time tokens $\mathbf{X}_{\text{time}}$ and the principal word embeddings $\mathbf{\hat{D}}$ through a Multi-head Cross-Attention layer. This lets us align the time tokens and word embeddings in the forecasting model’s embedding space to obtain the aligned cross-modal text tokens, i.e.,
(4) $\begin{gathered}\mathbf{X}_{\text{text}}=\text{Softmax}\left(\frac{\mathbf{Q}\mathbf{K}^{\top}}{\sqrt{C}}\right)\mathbf{V},\\ \text{where}\quad\mathbf{Q}=\mathbf{X}_{\text{time}}\mathbf{W}_{q},\quad\mathbf{K}=\hat{\mathbf{D}}\mathbf{W}_{k},\quad\mathbf{V}=\hat{\mathbf{D}}\mathbf{W}_{v}.\end{gathered}$

### v2

@@
Abstract
While Large Language Models have been used to produce interpretable stock forecasts, they mainly focus on analyzing textual reports but not historical price data, also known as Technical Analysis. This task is challenging as it switches between domains: the stock price inputs and outputs lie in the time-series domain, while the reasoning step should be in natural language. In this work, we introduce Verbal Technical Analysis (VTA), a novel framework that combine verbal and latent reasoning to produce stock time-series forecasts that are both accurate and interpretable. To reason over time-series, we convert stock price data into textual annotations and optimize the reasoning trace using an inverse Mean Squared Error (MSE) reward objective. To produce time-series outputs from textual reasoning, we condition the outputs of a time-series backbone model on the reasoning-based attributes. Experiments on stock datasets across U.S., Chinese, and European markets show that VTA achieves state-of-the-art forecasting accuracy, while the reasoning traces also perform well on evaluation metrics judged by industry experts. Our code is available at: https://github.com/chen-jan/VTA.
footnotetext: * Equal contribution. † Corresponding author.
1 Introduction
With the advent of Large Language Models (LLMs), an increasingly popular application is in financial analysis (Wu et al., 2023; Xie et al., 2023). This spans a wide range of tasks, including financial question answering (FinQA) (Liu et al., 2025b; Qian et al., 2025), investment decision-making (Yu et al., 2025; 2024), and market forecasting (Yu et al., 2023; Koa et al., 2024). Majority of existing approaches primarily utilize the strong natural-language capabilities of LLMs to analyze financial reports or do sentiment analysis on social texts (see Table 1), but neglects interpretable analysis on stock price data, which arguably contain useful information for financial practitioners.
The current solutions from the general time-series domain are not yet sufficient for this task. Existing studies on time-series reasoning (Merrill et al., 2024; Chow et al., 2024) consistently report that LLMs struggle to reason over raw time-series inputs. Meanwhile, time-series LLMs (Jin et al., 2024; Liu et al., 2025a) often rely on reprogramming the embedding space, which produces time-series outputs but sacrifices verbal reasoning ability, which is an essential requirement for interpretable financial analysis. The closest effort is TimeCAP (Lee et al., 2025), which generates explanations by contextualizing the series with auxiliary information. However, its reasoning trace is derived from external data, and it produces classification label forecasts rather than full time-series trajectories.
Unlike other time-series data, financial time-series contains intrinsic interpretable signals which are widely studied by experts, known as Technical Analysis (Kirkpatrick II and Dahlquist, 2010). We use these signals to verbally analyze financial time-series and produce interpretable stock forecasts.
Table 1: Comparison of relevant works. Our work contributes a novel explainable financial signal for practitioners and produces some insights into how time-series forecasting can be made interpretable.
@@ 1 Introduction
Fin-R1 (Liu et al., 2025b), Fino1 (Qian et al., 2025) Financial Question Answering Financial Reports Textual Answers
FinMem (Yu et al., 2025), FinCon (Yu et al., 2024) Investment Decision-Making Text, Tabular, Audio Binary (Buy/Sell)
GPT-4 (Yu et al., 2023), SEP (Koa et al., 2024) Market Direction Forecasting News, Social Texts Binary or Quantile-Based
@@
Time-LLM (Jin et al., 2024), CALF (Liu et al., 2025a) Time-Series Forecasting General Time-Series Time-Series
TimeCAP (Lee et al., 2025) Time-Series Reasoning Time-Series + Auxiliary Classification + Reasoning
@@
Verbal Technical Analysis (VTA)
Time-Series Reasoning
Financial Time-Series Time-Series + Reasoning
@@
The use of LLMs for time-series reasoning is hindered by three main challenges. Firstly, current LLMs have limited capabilities in time-series forecasts. Some works have tackled this by modifying the embedding space to produce time-series outputs (Jin et al., 2024; Liu et al., 2025a), but this comes at the cost of interpretability as the LLM loses its natural language capability. Secondly, on a higher level, current LLMs are not known to have the ability to do verbal reasoning on time-series to produce accurate forecasts (Merrill et al., 2024; Chow et al., 2024). This involves understanding how to best analyze the predictive signals in the time-series data in an unsupervised manner. Thirdly, the reasoning trace of the LLM further needs to be converted into time-series output to produce useful stock forecasts. LLMs are typically fine-tuned on next-token predictions (Radford et al., 2019), and using direct time-series outputs would not produce the best forecasts, which we verify empirically.
@@
To address these problems, we present three key contributions. Firstly, we propose our Verbal Technical Analysis (VTA) framework, which combines a backbone time-series model (which we termed as “latent thinking”) with a reasoning LLM (termed as “verbal reasoning”) to produce interpretable stock time-series forecasts. This framework combines the strong pattern processing ability of state-of-the-art time-series models and the strong reasoning ability of LLMs to produce forecasts that are both accurate and interpretable. Secondly, for reasoning over time-series, the stock time-series data is converted into textual annotations (Lin et al., 2024) as inputs to the LLM. The reasoning trace is then optimized through a modified Group Relative Policy Optimization (GRPO) objective (Shao et al., 2024) that uses an inverse Mean Squared Error (MSE) reward scoring, which we termed as Time-GRPO. Thirdly, to produce time-series outputs from the reasoning traces, we condition (Ho and Salimans, 2022) the generated outputs from the time-series model on the reasoning-based attributes.
@@
To demonstrate the effectiveness of VTA, we perform extensive experiments across established stock baselines (Xu and Cohen, 2018) and additional stock data across the U.S., Chinese, and European markets. We show that our model forecasts achieve state-of-the-art results in prediction accuracy, while also being interpretable. In addition, evaluation by industry experts show that the reasoning traces score high across various evaluation metrics from literature. Finally, to justify the practical capability of the model, we form Markowitz portfolios across the prediction length and show that the portfolios formed from VTA forecasts can also perform well on investment metrics.
2 Related Works
@@
Financial Large Language Models. The rise of Large Language Models (LLMs) has spurred a growing body of research on their application in finance. The earliest works focus on developing general-purpose financial LLMs, such as BloombergGPT (Wu et al., 2023) and FinMA (Xie et al., 2023), by finetuning on a large set of financial corpora across multiple downstream tasks. Later works began to tackle the specific challenges of LLMs in finance. For example, works on financial question-answering (FinQA) (Liu et al., 2025b; Qian et al., 2025) focus on teaching LLMs to analyze financial reports, which requires the ability to read structured financial tables and extract insights from complex documents (Zhu et al., 2021). LLMs for investment decision-making (Yu et al., 2025; 2024) typically utilize multi-agent systems to handle different parts of an investment decision, including but not limited to document analysis, memory, risk control, etc. LLMs for financial forecasting (Yu et al., 2023; Koa et al., 2024) seeks to predict the direction in which the market will go. Typically, these works analyze textual sources in order to understand the sentiment or financial health of a company. Our work positions itself in this field by learning to produce interpretable forward-looking signals from financial time-series data, which could benefit all applications above.
@@
Time-Series Large Language Models. At the same time, there is also a growing body of research in utilizing LLMs in the time-series domain (Kong et al., 2025). LLMs for time-series (Jin et al., 2024; Liu et al., 2025a) leverage the large scale parameters and robust pattern recognition of LLMs by fine-tuning them for time-series forecasting tasks. However, these approaches typically modify the embedding space of the LLMs, making them lose their original language reasoning capabilities. Some works have also explored the ability of LLMs to reason over time-series data. It was found that language models are “remarkably bad” at zero-shot time-series reasoning (Merrill et al., 2024), whereas fine-tuning them using a latent encoder (Chow et al., 2024) shows some promising early results in reasoning over time-series for captioning. The closest work on time-series reasoning comes from TimeCAP (Lee et al., 2025), which perform forecasting by contextualizing the input time-series with auxiliary time-series information. It produces explanations by searching for similar historical contexts, and produces classification label forecasts, in textual form. Our work builds on this line of research by exploiting the predictive signals within financial time-series to produce interpretable time-series forecasts, providing some insights on the capabilities of LLMs in reasoning over time-series data.
3 Verbal Technical Analysis
@@
Figure 1: The Verbal Technical Analysis (VTA) framework. We first teach an LLM to reason over time-series data. The reasoning outputs are used to condition a time-series forecasting model, to produce forecasts with similar attributes. This results in forecasts with interpretable reasoning traces.
3.1 Problem Formulation
We consider the task of forecasting short-term future stock prices, based on a historical window of $T$ trading days. Let $\mathbf{X}=\{\mathbf{x}_{t-T+1},\mathbf{x}_{t-T+2},\cdots,\mathbf{x}_{t}\}$, where the input vector consists of the open price, high price, low price, volume traded, closing price and adjusted closing price, i.e., $\mathbf{x}_{t}=\left[o_{t},h_{t},l_{t},v_{t},c_{t},p_{t}\right]$. We aim to generate an output $\mathbf{Y}=\{\mathbf{v},\mathbf{y}\}$, which consists of the verbal reasoning trace $\mathbf{v}$ and the price trajectory over the next $T^{\prime}$ trading days $\mathbf{y}=\{p_{t+1},p_{t+2},\cdots,p_{t+T^{\prime}}\}$.
3.2 Time-Series Reasoning
To teach an LLM to reason over time-series inputs, we use a textual annotator to extract useful interpretable signals for forecasting. The LLM uses these indicators to reason over the time-series to make forecasts without any supervision data. This is achieved through our proposed Time-Series Group Relative Policy Optimization (Time-GRPO) method, which uses a multi-stage reinforcement learning (RL) pipeline, together with a modified GRPO objective (Shao et al., 2024).
The time-series input is first converted into textual annotations (Lin et al., 2024), which consist of its statistics information (Jin et al., 2024) (e.g., its mean, minimum and maximum values) and financial technical indicators (Murphy, 1999) (e.g., moving averages, momentum, etc.). Formally, we have:
$\mathbf{X^{\prime}}=\mathbf{f}(\mathbf{X}),$ (1)
where $\mathbf{f}$ contains the annotation functions and $\mathbf{X^{\prime}}$ are the annotated values. A full list of the financial technical indicators used, with their descriptions and calculations, are provided in Appendix B.
@@ 3.2 Time-Series Reasoning
$\mathcal{L}_{\text{time-grpo}}(\theta)$ $=\mathbb{E}_{\mathbf{q}\sim\mathbf{\mathcal{Q}},\{\mathbf{o_{i}}\}_{i=1}^{G}\sim\pi_{\theta_{\text{old}}}(\mathbf{O}|\mathbf{q})}$
$\frac{1}{G}\sum_{i=1}^{G}\left(\min\left(\frac{\pi_{\theta}(\mathbf{o_{i}}|\mathbf{q})}{\pi_{\theta_{\text{old}}}(\mathbf{o_{i}}|\mathbf{q})}A_{i},\,\text{clip}\left(\frac{\pi_{\theta}(\mathbf{o_{i}}|\mathbf{q})}{\pi_{\theta_{\text{old}}}(\mathbf{o_{i}}|\mathbf{q})},1-\epsilon,1+\epsilon\right)A_{i}\right)-\beta\mathbb{D}_{\text{KL}}(\pi_{\theta}\|\pi_{\text{ref}})\right),$ (2)
$\quad\quad\quad\quad\quad\mathbb{D}_{\text{KL}}(\pi_{\theta}\|\pi_{\text{ref}})=\frac{\pi_{\text{ref}}(\mathbf{o_{i}}|\mathbf{q})}{\pi_{\theta}(\mathbf{o_{i}}|\mathbf{q})}-\log\frac{\pi_{\text{ref}}(\mathbf{o_{i}}|\mathbf{q})}{\pi_{\theta}(\mathbf{o_{i}}|\mathbf{q})}-1,$ (3)
where $\epsilon$ and $\beta$ are hyper-parameters. $A_{i}$ denotes the advantage of the LLM policy, which is derived from a set of rewards $\{r_{1},r_{2},\cdots,r_{G}\}$ that are associated with outputs $\mathbf{O}$ produced in each group:
$A_{i}=\frac{r_{i}-\text{mean}(\{r_{1},r_{2},\cdots,r_{G}\})}{\text{std}(\{r_{1},r_{2},\cdots,r_{G}\})}.$ (3)
We utilize the format reward that was used in previous works (Guo et al., 2025), that enforces the model to always employ a thinking process that is between <think><think> and </think></think> tags.
Ideally, the generated reasoning trace should also maximize the expected accuracy of the time-series forecasts. This is achieved by utilizing the Mean-Squared Error (MSE) score as an additional reward:
$r_{\text{MSE}}(\theta)=1/\left(\lambda\cdot\left\|\hat{\mathbf{y}}_{\theta}-\mathbf{y}\right\|_{2}^{2}\right),$ (4)
@@
Training Pipeline. Following established practices in LLM fine-tuning literature (Guo et al., 2025; Ouyang et al., 2022; Lu et al., 2024; Chow et al., 2024), Time-GRPO utilizes a multi-stage pipeline to fine-tune the time-series reasoning LLM.
[…]
3.3 Time-Series Forecasting
To perform time-series forecasting, we employ LLM-based time-series models. Works have shown that the powerful contextual modeling capabilities of LLMs can be effectively adapted for time-series forecasting tasks (Zhou et al., 2023; Chang et al., 2023; Cao et al., 2023). A key technique is to align the time-series and language distributions (Sun et al., 2023; Jin et al., 2024) such that the model is able to understand the context of time-series data (e.g., up, down, steady, etc.). For our backbone model, we repurpose an LLM for cross-modal fine-tuning (Liu et al., 2025a).
For this step, we first pass the time-series input $\mathbf{X}$ through an embedding layer, followed by a multi-head attention layer, to obtain the projected time tokens $\mathbf{X}_{\text{time}}$. Next, it is observed that similar words are usually close to each other in the LLM embedding space, and for non-text based tasks, it is sufficient to keep cluster centers to reduce training costs (Sun et al., 2023; Liu et al., 2025a). To do this, we perform Principal Component Analysis (PCA) to retrieve the principal word embeddings $\mathbf{\hat{D}}$. Following this, we then pass the projected time tokens $\mathbf{X}_{\text{time}}$ and the principal word embeddings $\mathbf{\hat{D}}$ through a Multi-head Cross-Attention layer. This lets us align the time tokens and word embeddings in the forecasting model’s embedding space to obtain the aligned cross-modal text tokens, i.e.,
$\begin{gathered}\mathbf{X}_{\text{text}}=\text{Softmax}\left(\frac{\mathbf{Q}\mathbf{K}^{\top}}{\sqrt{C}}\right)\mathbf{V},\\ \text{where}\quad\mathbf{Q}=\mathbf{X}_{\text{time}}\mathbf{W}_{q},\quad\mathbf{K}=\hat{\mathbf{D}}\mathbf{W}_{k},\quad\mathbf{V}=\hat{\mathbf{D}}\mathbf{W}_{v}.\end{gathered}$ (5)